OpenStudy (anonymous):

the lenght of a median of a trapezoid is 7 units the lenght of a longer base is 21 units how long is the shorter base

OpenStudy (anonymous):

how can i solve it...?

OpenStudy (callisto):

I think using the formula @waterineyes gave you in the previous post: \[\text{Length Of Median} = \frac{1}{2} \times (\text{Sum Of Length Of Bases)}\]

OpenStudy (anonymous):

Here you are given : \[7 = \frac{1}{2} \times (21 + x)\] Find x here, that is your Length of Shorter Base..

OpenStudy (anonymous):

waterineyes: m=b1+b2/2 can be this formula